# Cavity-Altered Superconductivity

### v3

@@
Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment? Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8, 9], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [10, 11, 12, 13, 14, 15]. Here, we devised and implemented a novel platform to realize cavity-altered materials. Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and prominent mode confinement [16, 17, 18, 19, 20, 21, 22]. We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET). The frequencies of infrared (IR) hyperbolic modes of hBN [23, 24] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [25, 26]. Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm the presence of resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET. Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface. Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the superfluid suppression. These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state. This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground state properties.
@@
Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice. The emergent effects born out of such strong interactions are abundant and drive the formation of varied electronic and magnetic phases. At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter. A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring the quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8, 9]. Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase in electron-phonon coupling via phonon polaritons [1, 21, 7]; Amperean pairing directly mediated through dark-cavity photons [27, 28]; and the driven-cavity extension of the Eliashberg effect [29]. Experiments with dark cavities revealed modified transport properties in the integer and fractional quantum Hall states of a 2D electron gas [10, 13, 15] as well as cavity-assisted thermal control of the metal-to-insulator transition in charge density wave systems [11].
@@
We focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings. We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts below the transition temperature =11.5 K [30, 31] (Supplementary Information Sections 1,2). We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN). Hyperbolicity occurs when the permittivity has opposite signs along different axes [23, 24] and results in a highly enhanced photonic density of states [32, 33, 34]. Notably, hyperbolic modes of hBN overlap with the frequency of the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [35] and the two modes hybridize at the hBN/-ET interface. Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode of -ET that is implicated in superconductivity [25, 26]. In this work, we examine how superconductivity is altered at the hBN/-ET interface, using advanced scanning probe-based techniques that effectively analyze the electrodynamics of buried interfaces.
Meissner force probe of resonant and non-resonant interfaces
@@
Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [37, 38, 39]. The force results from the supercurrent screening the stray magnetic field of the tip. Notably, the Meissner force is not affected when non-magnetic materials are inserted between the tip and the superconductor, and the data in Fig. 1(a,b) attest that MFM probes superconductors encapsulated by insulating layers. We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [40]. Analysis of allows the extraction of the superfluid density [38] (Methods and Supplementary Information Section 4).
@@
We examine the impact of hBN on superconductivity in a representative heterostructure composed of a 60-nm-thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick). We report measured as a function of , the separation between the tip and the -ET surface, in Fig. 1a. On a region of the -ET crystal unobscured by hBN (green), is negative at K, indicating a repulsive force. The magnitude of grows with reduced tip-sample surface separation. The signal collected above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading. As a control experiment, we placed a 55-nm-thick RuCl3 microcrystal on top of -ET. RuCl3 is an insulator with a static permittivity similar to that of hBN [41, 36]. However, optical phonons of RuCl3 all occur at much lower frequencies below 350 cm-1 [36] and therefore do not resonantly couple to the C=C stretching mode of -ET. Fig. 1c schematically depicts the permittivities of hBN, -ET, and RuCl3. Comparing measurements above RuCl3/-ET (Fig. 1b, magenta curve) and bare -ET (green curve - same as in Fig. 1a), we observe a much weaker (less than %) effect than over hBN/-ET. The stark contrast between the RuCl3/-ET and hBN/-ET interfaces indicates that the suppression of superfluid density under hBN is likely prompted by resonant coupling across the hBN/-ET interface. We remark that the RuCl3 control experiment rules out a prominent role for work function disparity [42] in the suppression of the superfluid density in Fig. 1 and similarly suggests that strain is unlikely to be significant. We perform a second control measurement on a heterostructure of 8-nm hBN residing on a 28.5-nm-thick microcrystal of (BSCCO) with K. Crystals of BSCCO host phonons below 650 cm-1 [43], far below the hyperbolic modes in hBN (Supplementary Information Section 5). MFM data plotted in Fig. 1b (inset) for hBN/BSCCO (blue) are indistinguishable from the data for bare BSCCO (dark gray). Thus, the significant suppression of the superfluid density is specific to the resonant nature of the hBN/-ET interface.
@@
Next, we investigate resonant mode coupling across the hBN/-ET interface. The hBN crystal hosts hyperbolic phonon polaritons (HPhPs) that exhibit light-like dispersion near the transverse optical (TO) frequency and phonon-like dispersion near the longitudinal optical (LO) frequency [16]. Here, TO and LO refer specifically to the in-plane modes. We examine the HPhP dispersion upon interfacing with -ET through the calculated imaginary part of the p-polarized Fresnel reflection , as shown in Fig. 2a (Supplementary Information Section 6). For free-standing hBN, the HPhP dispersion has a series of branches between the TO and LO phonon frequencies, each with a continuous dispersion [23]. By contrast, for an hBN/-ET heterostructure, the HPhP branches are interrupted near the C=C stretching frequency. Avoided crossings are visible, yet the gaps in the HPhP dispersion are not fully opened due to broadening caused by dielectric loss associated with the in-plane metallic response of -ET [45, 46].
@@
To experimentally examine the mode coupling at the hBN/-ET interface, we used scattering-type scanning near-field optical microscopy (s-SNOM) [47]. Specifically, we evaluated the hBN HPhP dispersion as these modes couple to the C=C stretching mode of -ET. We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser. The strong electric field with high momentum achieved near the tip apex launches polariton waves. We performed nano-IR measurements at K on a region of an hBN/-ET heterostructure with 32-nm hBN on 100-nm -ET. Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images: representative results are shown in Fig. 2b for cm-1. Line profiles from s-SNOM collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig. 2c). In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with the frequency [48]. However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the solid line tracking the fifth peak in Fig. 2c. These nano-IR data attest to mode coupling across the hBN--ET interface.
@@
To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HPhP and the C=C stretching mode. First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP’s electric field with the C=C stretching mode. The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig. 2d and Supplementary Information Section 2). It is important to note that the HPhP field has an out-of-plane component that can couple to the C=C mode (Fig. 2d). The combined contributions of many HPhP branches outspreading to high momenta make this field extend far beyond the hBN surface. We therefore calculate the C=C amplitude in the presence of an oscillating out-of-plane electric field with a frequency spectrum spanning the C=C frequency (fundamental frequency cm-1). Langevin molecular dynamics analysis carried out at K (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude and splits the peak in Fig. 2e. We therefore conclude that vacuum fluctuations—specifically, zero-point phonon motion—mediate the coupling between hyperbolic phonon polaritons and molecular vibrations without the need for external photons, thus enabling interaction in the dark.
Suppression of superfluid density at the hBN/-ET interface
We now validate the association of reduced (Fig. 1a) with suppressed superfluid density by studying the temperature dependence. Over bare -ET, the large repulsive is rooted in the Meissner effect (Supplementary Information Section 4), it decreases with increasing temperature (Fig. 3a) and vanishes near . Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effects at the heterostructure interfaces by performing constant-height scans with the tip hovering above the -ET surface in a 1616-m2 region that includes both hBN and RuCl3 (Fig. 3b-d). To facilitate comparison between temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET. At 2 K (panel b), pN/m over hBN, indicating a weaker Meissner force. As with the curves in Fig. 1(a,b), the effect over RuCl3 is negligible. Warming to 12 K, just above , the image contrast disappears, giving everywhere. The vanishing near confirms that the weakened observed at 2 K is associated with a reduction in the superfluid density. Further evaluation and additional control measurements are described in Supplementary Information Section 8.
We observe suppression of superfluid density under hBN in multiple hBN/-ET heterostructures. The magnitude of suppression varies spatially (Supplementary Information Section 9); we focus on regions with the strongest effect. The results are summarized in Fig. 4. To model the Meissner force analytically [37], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth and evaluate the effective superfluid density under hBN, . In Fig. 4, we compile MFM data for several devices with hBN cavities and RuCl3 dielectric layers of varied thicknesses. We present data in the form of , where is the superfluid density in bare -ET. In all heterostructures, we observe at least 50% suppression of the superfluid density prompted by the hBN cavity. This result holds over a large variation in the hBN thickness, ranging from 25 nm to 110 nm. Accurate conversion between and superfluid density would require knowledge of the depth profile (Supplementary Information Section 4). While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate among the possible depth profiles that all lead to similar Meissner forces. However, the magnitude of the observed Meissner force suppression implies that -ET is impacted by the cavity to a significant depth. The impacted depth appears to be of the same order of magnitude as the polaritonic wavelength (Supplementary Information Section 10).
@@
The data in Figs. 1-4 demonstrate that cavity electrodynamics plays a pivotal role in suppressing the superfluid density – a hallmark equilibrium property of superconductivity. These experiments establish a precedent for successfully altering a material’s ground state thermodynamic property through cavity engineering. Our specific implementation of the resonant structure reveals two notable features of hyperbolic cavity control. First, hyperbolic modes entail both in-plane and out-of-plane components of the fluctuating electromagnetic fields. The out-of-plane component is crucial for coupling to the out-of-plane dipole moment of the C=C stretching mode in the -ET superconductor under investigation. Second, the multiple branches of hyperbolic modes span a broad momentum range, enabling continuous interaction with the weakly dispersing C=C resonance across the entire phase space, as shown in Fig. 2a. Future experiments will be needed to evaluate the spatial extent of the mode coupling away from the hyperbolic interface; our earlier data for a similar resonant hBN/MoO3 heterostructure suggest that the relevant length scale may exceed nm [49].
Our data and modeling suggest a plausible pathway for cavity enhancement of superconducting properties. Indeed, our simulations indicate that the hBN cavity subdues the C=C stretching mode fluctuations prompting the stark suppression of superfluid density. By the same token, boosting fluctuations through mode softening in the cavity [1] may increase the superfluid density and possibly in -ET. Furthermore, the concept of employing readily reconfigurable vdW hyperbolic cavities for resonant manipulations of material properties is rather general. We remind the reader that vdW crystals host hyperbolic modes of phononic, plasmonic and exitonic origin [50] spanning the range from THz to visible, thus offering limitless options for hyper-cavity engineering.
@@
- [32] Poddubny, A., Iorsh, I., Belov, P. & Kivshar, Y. Hyperbolic metamaterials. Nat. Photonics 7, 948–957 (2013).
@@
- [47] Chen, X. et al. Modern scattering-type scanning near-field optical microscopy for advanced material research. Adv. Mater. 31, 1804774 (2019).
- [48] Bylinkin, A. et al. Real-space observation of vibrational strong coupling between propagating phonon polaritons and organic molecules. Nat. Photonics 15, 197–202 (2021).
@@
- [50] Basov, D. et al. Polaritonic quantum matter. Nanophotonics (2025).

### v4

@@
Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment? Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [15, 16, 17, 18, 19, 20, 21]. Here, we devised and implemented a novel platform to realize cavity-altered materials. Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and prominent mode confinement [22, 23, 24, 25, 26]. We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET). The frequencies of infrared (IR) hyperbolic modes of hBN [27, 28] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [29]. Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm the presence of resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET. Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface. Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the pronounced superfluid suppression. These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state. Our work highlights the potential of dark cavities devoid of external photons for engineering electronic ground state properties of complex quantum materials.
@@
Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice. The emergent effects born out of such strong interactions are abundant and drive the formation of varied electronic and magnetic phases. At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter. A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring the quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Experiments with dark cavities revealed modified transport properties in the integer and fractional quantum Hall states of a 2D electron gas [15, 19, 21] as well as cavity-assisted thermal control of the metal-to-insulator transition in charge density wave systems [17]. Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase in electron-phonon coupling via phonon polaritons [1, 10, 11], Amperean pairing directly mediated through dark-cavity photons [30, 31] as well as the driven-cavity extension of the Eliashberg effect [32].
@@
Here, we have chosen to focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings. We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts below the transition temperature =11.5 K [33, 34] (Supplementary Information Sections 1,2). We harness an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN). Hyperbolicity occurs when the permittivity has opposite signs along different axes [27, 28] and results in a highly enhanced photonic density of states [35, 36, 37]. Notably, hyperbolic modes (HMs) of hBN overlap with the frequency of the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [38] and the two modes hybridize at the hBN/-ET interface. Effectively, the hBN slab acts as an electromagnetic cavity (Supplementary Information Section 3) resonantly tuned to the C=C stretching mode of -ET that is implicated in superconductivity [29]. In this work, we examine how superconductivity is altered at the hBN/-ET interface, using advanced scanning probe-based techniques adept at the analysis of electrodynamics in buried interfaces.
Meissner effect at resonant and non-resonant interfaces
@@
Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [41, 42, 43, 44]. The force results from the supercurrent screening the stray magnetic field of the tip. Notably, the Meissner force is not affected when non-magnetic materials are inserted between the tip and the superconductor, and the data in Fig. 1(a,b) attest that MFM probes superconductors encapsulated by insulating layers. We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [45]. Analysis of allows the extraction of the superfluid density [42] (Methods and Supplementary Information Section 5).
@@
We examine the impact of hBN on superconductivity in a representative heterostructure composed of a 60-nm-thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick). We report measured as a function of , the separation between the tip and the -ET surface, in Fig. 1a. On a region of the -ET crystal unobscured by hBN (green), is negative at K, indicating a repulsive force. The magnitude of grows with reduced tip-sample surface separation. The signal collected above hBN (blue) shows a weakened , highlighted by the gray shading. As a control experiment, we placed a 55-nm-thick RuCl3 microcrystal on top of the same -ET sample. RuCl3 is an insulator with a static permittivity similar to that of hBN [46, 40]. However, optical phonons of RuCl3 all occur at much lower frequencies below 350 cm-1 [40] and therefore do not resonantly couple to the C=C stretching mode of -ET. Fig. 1c schematically depicts the permittivities of hBN, -ET, and RuCl3. Comparing measurements above RuCl3/-ET (Fig. 1b, magenta curve) and bare -ET (green curve - same as in Fig. 1a), we observe a much weaker (less than %) effect than over hBN/-ET.
The stark contrast between the RuCl3/-ET and hBN/-ET interfaces indicates that the suppression of superfluid density under hBN is likely prompted by resonant coupling across the hBN/-ET interface. We remark that the RuCl3 control experiment rules out a prominent role of charge transfer due to work function disparity [47] in the suppression of the superfluid density in Fig. 1 (Supplementary Information Section 14) and similarly suggests that strain is unlikely to be significant. We perform a second control measurement on a heterostructure of 8-nm hBN residing on a 28.5-nm-thick microcrystal of (BSCCO) with K. Crystals of BSCCO host phonons below 650 cm-1 [48], far below the HMs in hBN. MFM data plotted in Fig. 1b (inset) for hBN/BSCCO (blue) are indistinguishable from the data for bare BSCCO (dark gray - Supplementary Information Section 6). Thus, the significant suppression of the superfluid density is specific to the resonant nature of the hBN/-ET interface.
Suppression of superfluid density at the hBN/-ET interface
We now validate the association of reduced (Fig. 1a) with suppressed superfluid density by studying the temperature dependence. Over bare -ET, the large repulsive arises from the Meissner effect (Supplementary Information Section 5), decreases with increasing temperature (Fig. 2a), and vanishes near . Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effects at the heterostructure interfaces. We performed constant-height scans with the tip hovering above the -ET surface in a 1616 m2 region that includes both hBN and RuCl3 (Fig. 2b,c). To facilitate comparison between data collected at different temperatures, we focus on the differential signal , where the latter term stands for over bare -ET. At 2 K (Fig. 2b), pN/m over hBN, indicating suppressed superfluid density. Over RuCl3, is visibly much smaller. More quantitatively, in Fig. 2d we display the histogram corresponding to the image of the signals. We witness distinct distributions representing bare -ET, RuCl3/-ET, and hBN/-ET. The hBN/-ET peak is prominently separated from the other two, indicating the strong effect of hBN on the superfluid density of -ET. Supplementary Information Sections 7 and 8 discuss additional effects, including possible scenarios for the minor difference between the MFM data registered at the RuCl3/-ET interface and the bare -ET. Upon warming to 12 K, just above , the image contrast disappears, giving everywhere (Fig 2c). The vanishing near confirms that the weakened signal observed at 2 K is associated with a reduction in the superfluid density.
We observe suppression of superfluid density under hBN in multiple hBN/-ET heterostructures. The magnitude of suppression varies spatially (Supplementary Information Section 8); we focus on regions with the strongest effect. The results are summarized in Fig. 3. To model analytically [41], we make the grossly simplifying assumption that the suppression of the superfluid density is uniform in depth and evaluate the effective superfluid density under hBN, . In Fig. 3, we compile MFM data for several devices with hBN cavities and RuCl3 dielectric layers of varied thicknesses. We present data in the form of , where is the superfluid density in bare -ET. In all heterostructures, we observe at least 50% suppression of the superfluid density prompted by the hBN cavity. This result holds over a large variation in the hBN thickness, ranging from 25 nm to 110 nm. Accurate conversion between and superfluid density would require knowledge of the depth profile (Supplementary Information Section 5). While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate among the possible depth profiles that all lead to similar . Nevertheless, the magnitude of the observed suppression of implies that superconductivity in -ET is impacted by the cavity to a significant depth (Supplementary Information Section 9).
@@
Next, we investigate resonant mode coupling across the hBN/-ET interface. The hBN crystal hosts HMs that exhibit light-like dispersion near the transverse optical (TO) frequency and phonon-like dispersion near the longitudinal optical (LO) frequency [22]. Here, TO and LO refer specifically to the in-plane modes. In practice, the HMs manifest as hyperbolic phonon-polaritons (HPhPs) that can be visualized in nano-optical experiments as originally done in Ref. [27]. We examine the HPhP dispersion of hBN upon interfacing with -ET through the calculated imaginary part of the p-polarized Fresnel reflection , as shown in Fig. 4a (Supplementary Information Section 10). For free-standing hBN, the HPhP dispersion displays a series of branches between the TO and LO phonon frequencies, each with a continuous dispersion [27]. By contrast, for an hBN/-ET heterostructure, the HPhP branches are interrupted near the C=C stretching frequency (light blue box in Fig. 4a). Avoided crossings are visible, yet the gaps in the HPhP dispersion are not fully developed due to broadening caused by dielectric loss associated with the in-plane response of -ET [50, 51] (Supplementary Information Section 10).
@@
To experimentally examine the mode coupling at the hBN/-ET interface, we used scattering-type scanning near-field optical microscopy (s-SNOM) [52]. Specifically, we evaluated the HPhP dispersion of hBN as its modes couple to the C=C stretching mode of -ET. We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser. The strong electric field with high momentum achieved near the tip apex launches polaritonic waves. We performed nano-IR measurements at K on a region of an hBN/-ET heterostructure with 32-nm hBN on 100-nm -ET. Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images: representative results are shown in Fig. 4b for cm-1. Line profiles from s-SNOM collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig. 4c). In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with the frequency [53]. However, in our data on -ET-supported hBN, the fringe evolution exhibits an evident kink near the C=C frequency, indicated by the solid line tracking the fifth peak in Fig. 4c. These nano-IR data attest to mode coupling across the hBN--ET interface
@@
To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HMs and the C=C stretching mode. First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the electric field of the HMs with the C=C stretching mode. The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig. 4d and Supplementary Information Section 2). It is important to note that the electric field of the HM has an out-of-plane component that can couple to the C=C mode (Fig. 4d). The combined contributions of many HM branches reaching high momenta make this field extend far beyond the hBN surface (Supplementary Information Section 3). We therefore calculate the C=C amplitude in the presence of an oscillating out-of-plane electric field with a frequency spectrum spanning the C=C frequency (fundamental frequency cm-1). Langevin molecular dynamics analysis carried out at K (Supplementary Information Section 11) shows that the out-of-plane electric field induced by the zero-point HM fluctuations reduces the C=C amplitude and splits the peak in Fig. 4e. We therefore conclude that zero-point HM motion mediates the coupling between HMs and molecular vibrations without the need for external photons.
@@
The data in Figs. 1-4 demonstrate that cavity electrodynamics plays a pivotal role in suppressing the superfluid density – a hallmark equilibrium attribute of superconductivity. These experiments also establish a precedent for successfully altering a material’s ground state thermodynamic property through cavity engineering. Our specific implementation of the resonant structure exploits two notable features of hyperbolic cavity control. First, HMs entail both in-plane and out-of-plane components of the fluctuating electromagnetic fields. The out-of-plane component is crucial for coupling to the primarily out-of-plane dipole moment of the C=C stretching mode in the -ET superconductor under investigation. Second, the multiple branches of the HMs enable interaction with the weakly dispersing C=C resonance at multiple intersections over a broad range of momenta, as shown in Fig. 4a. Future experiments need to evaluate the spatial extent of the mode coupling away from the hyperbolic interface; our earlier data for a similar resonant hBN/MoO3 heterostructure suggest that the relevant length scale may exceed nm [54].
Our data and modeling suggest a plausible pathway for cavity enhancement of superconductivity. As an illustrative example, we consider the scenario of an optical phonon in a superconductor coupled to a hBN cavity within the framework of Bardeen, Cooper and Schrieffer (BCS) theory (Supplementary Information Section 13). Electron-phonon coupling, quadratic in the phonon coordinate [1, 55], results in a suppression of the superfluid density, while linear coupling results in enhancement. This analysis suggests that conventional superconductors may exhibit increased superfluid density through interactions with HMs. Here we pause to caution the reader: -ET studied in our work is an unconventional superconductor [56], and BCS-derived inferences can only guide our intuition. A rigorous description of the physics involved in cavity-coupled -ET awaits further investigation. We conclude by highlighting that the concept of employing readily reconfigurable vdW hyperbolic cavities for resonant manipulation of material properties is rather general. HMs of phononic, plasmonic and excitonic origin are hosted in many vdW materials including insulators, semiconductors as well as a variety of emerging materials [26]. These modes span the range from THz to visible [57], thus offering limitless options for hyper-cavity engineering.
@@
- [13] Fassioli, F., Faist, J., Eckstein, M. & Fausti, D. Controlling radiative heat flow through cavity electrodynamics. Phys. Rev. B 111, 165425 (2025).
- [14] Riolo, R. et al. Tuning Fermi liquids with polaritonic cavities. PNAS 122, e2407995122 (2025).
@@
- [16] Appugliese, F. et al. Breakdown of topological protection by cavity vacuum fields in the integer quantum Hall effect. Science 375, 1030–1034 (2022).
@@
- [35] Poddubny, A., Iorsh, I., Belov, P. & Kivshar, Y. Hyperbolic metamaterials. Nat. Photon. 7, 948–957 (2013).
@@
- [39] Giles, A. et al. Ultralow-loss polaritons in isotopically pure boron nitride. Nat. Mater. 17, 134–139 (2018).
@@
- [44] Grebenchuk, S. Y. et al. Crossover from ferromagnetic superconductor to superconducting ferromagnet in P-doped EuFe2(As1-xPx)2. Phys. Rev. B 102, 144501 (2020).
@@
- [52] Hillenbrand, R., Abate, Y., Liu, M., Chen, X. & Basov, D. N. Visible-to-THz near-field nanoscopy. Nat. Rev. Mater. 10, 285–310 (2025).
- [53] Bylinkin, A. et al. Real-space observation of vibrational strong coupling between propagating phonon polaritons and organic molecules. Nat. Photon. 15, 197–202 (2021).
@@
- [55] Kennes, D. M., Wilner, E. Y., Reichman, D. R. & Millis, A. J. Transient superconductivity from electronic squeezing of optically pumped phonons. Nat. Phys. 13, 479–483 (2017).
- [57] Ruta, F. L. et al. Good plasmons in a bad metal. Science 387, 786–791 (2025).